OpenStudy (anonymous):

A model rocket is launched with an initial upward velocity of 204/fts. The rocket's height h (in feet) after t seconds is given by the following. h=204t-16t^2 Find all values of t for which the rocket's height is 100 feet.

OpenStudy (anonymous):

Plug 100 into h and then solve the Quadratic equation.

OpenStudy (anonymous):

I did that and subtracted from both sides but then when I tried to divide the terms by -16 it didn't really work (I'm just going off of what I learned earlier)..

OpenStudy (anonymous):

You should end up with this equation \[-16t^2+204t-100=0\]

OpenStudy (anonymous):

I did but aren't you suppose to then divide -16 by the terms to then be able to use the quadratic formula?

OpenStudy (anonymous):

Then use the quadratic equation which is: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

Where a= -16, b= 204, and c=-100

OpenStudy (anonymous):

And no you don't have to divide both sides by -16. Once you have your equation in the proper format you just plug the coefficients into the equation and get your answer.

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

so far I have \[-204\pm \sqrt{\frac{ 35216 }{ -32 }}\]

OpenStudy (anonymous):

Good except the bottom should not be in the square root

OpenStudy (anonymous):

\[{\frac{ -204\pm \sqrt{35216} }{ -32}}\]

OpenStudy (anonymous):

Oh okay, I don't know what exactly to do after this step.

OpenStudy (anonymous):

Well You solve the above equation twice. Once with the plus or minus being a plus. And another time with it being a minus

OpenStudy (anonymous):

I got .51 and 12.24

OpenStudy (anonymous):

Yep correct those are your two values of t at the height of 100